# Approach to stationarity for the KPZ fixed point with boundaries

### v1

@@ Abstract
$\langle\mathrm{e}^{sh(x,\tau)}\rangle\simeq\mathrm{e}^{\tau\mu(s)+\nu_{x}(s)}\quad\text{when {\hbox{\tau\to\infty}}}\;.$→∞ . (3)
@@
A classical mapping of TASEP to an interface growth model, see figure 1, interprets empty (resp. occupied) sites as portions of interface going up (resp. down). This defines a bijection between configurations of the particles on the lattice and discrete height functions, written from occupation numbers $n_{i}$ (with $n_{i}=1$ if site $i$ is occupied and $n_{i}=0$ if site $i$ is empty) as
@@
$i=0,\ldots,L$, with the convention $H_{0}=0$. Movements of particles on the lattice then induce deposition of square blocks on the local minima of the interface, with $H_{i}\to H_{i}+1$ when a particle moves from site $i$ to site $i+1$ (with $i=0$ for particles entering at site $1$ and $i=L$ for particles leaving from site $L$). The resulting time dependent height function $H_{i}(t)$ is initialized at time $t=0$ as (5) with the occupation numbers of the initial configuration (which may be random). Additionally, the time-integrated current $Q_{i}(t)$, defined as the total number of particles that have moved from site $i$ to $i+1$ up to time $t$, is equal to
@@
In the whole maximal current phase $\alpha>1/2$, $\beta>1/2$ of TASEP [27], the large scale limit $L\to\infty$ of the discrete height on the time scale $t\sim L^{3/2}$ leads to the KPZ fixed point with infinite boundary slopes, and we set already $\alpha=\beta=1$ for simplicity in the following. Considering a smooth enough continuous function $h_{0}(x)$, $0\leq x\leq 1$, with $h_{0}(0)=0$, we choose an initial condition of the form
@@
The matrix elements of the generator $M_{i}(\gamma)$ are equal to those of the Markov operator of TASEP, multiplied by $\mathrm{e}^{\gamma}$ for all transitions where a particle moves from site $i$ to site $i+1$. One has in particular $M_{i}(\gamma)=V_{i}^{-1}(\gamma)M_{0}(\gamma)V_{i}(\gamma)$ with $V_{i}$ diagonal such that $\langle C|V_{i}|C\rangle=\mathrm{e}^{-\gamma H_{i}}$, and $H_{i}$ related to $C$ through (5). The row vector $\langle\Sigma|=(1,1,\ldots,1)$ in (10) leads to a summation over all possible final states, and $|\mathbb{P}_{0}\rangle$ is the column vector of initial probabilities.
@@
where $w_{1}(u)$ and $w_{2}(u)$ are associated with $C^{1}$ and $C^{2}$, and the minimum is taken over all $u\in[0,1]$. Similarly, $\langle C^{2}|U|C^{1}\rangle$ leads to conditioning the corresponding $w_{1}$ and $w_{2}$ on $w_{1}(u)\geq w_{2}(u)$ for $u\in[0,1]$ and $w_{1}(1)=w_{2}(1)$. Equivalently, $w(x)=(w_{1}(x)+w_{2}(x))/\sqrt{2}$ and $e(x)=(w_{1}(x)-w_{2}(x))/\sqrt{2}$ are respectively a Wiener process and a Brownian excursion, independent from each other, see the Supplemental Material [32]. Then, (11) combined with (6), (Approach to stationarity for the KPZ fixed point with boundaries), (13) and (14) finally leads to our first main result, valid for any $n\in\mathbb{N}$ and arbitrary initial condition $h_{0}$
@@
The average height $\langle h(x,\tau)\rangle$ has the form (4) with the known value $c_{1}=3/2$. From (18), the new correction to stationarity (see the Supplemental Material [32] for the computation of the required Brownian averages) is
@@
The $L+2$ solutions $w_{j}$, $j=0,\ldots,L+1$ of (22) lead to $y_{j}=y_{j}^{\pm}=w_{j}\pm w_{j}\sqrt{1-w_{j}^{-2}}$. Taking the principal branch for the square root, our numerics indicate that any eigenstate of $M_{i}(\gamma)$ can be generated by choosing Bethe roots $y_{j}=y_{j}^{\theta_{j}}$ with the signs $\theta_{j}=\pm$ such that $\prod_{j=0}^{L+1}\theta_{j}=(-1)^{L}$. The stationary eigenstate corresponds in particular to selecting $y_{j}=y_{j}^{-}$ for all $j$.
@@
and $\omega^{\rm full}=\omega^{\rm empty}-\frac{L}{2}\,\kappa\,\frac{\mathrm{d}B}{B}$. These conjectures were checked against numerical diagonalization of $M_{0}(\gamma)$ up to system size $L=7$, with $100$ digits of precision for all acceptable choices $y_{j}^{\pm}$ for a few generic values of $B$.
When $L\to\infty$ under KPZ scaling (15) for $\gamma$, setting $B/B_{*}=-\mathrm{e}^{v-1}$, with $B_{*}$ the finite non-zero branch point of the $y_{j}^{\pm}$, a residue computation leads for the stationary eigenstate to $\kappa_{\rm st}\simeq 2\chi^{\prime\prime}(v)/\sqrt{L}$ with $\chi$ as in (16) and $v<1$ the unique solution of $\chi^{\prime}(v)=s$ for $s\in(0,s_{*}]$ with $s_{*}=\chi^{\prime}(1)\approx 0.145931$. Noting that $\mathcal{N}(B)$ evaluated at the stationary eigenstate appears in (11), we finally obtain our second main result for infinite boundary slopes,
@@
for narrow wedge initial condition at the left boundary, using $\omega^{\rm empty}$ in (24). The same result is obtained for a narrow wedge at the right boundary, using $\omega^{\rm full}$, with an additional deterministic subtraction in (9), see the Supplemental Material [32]. For stationary initial condition $h_{0}=h_{\rm st}$ defined in (2), one finds instead from (23)
@@
Supplemental Material
We give in this supplemental material a more detailed derivation of the main results presented in the letter for the function $\nu_{x}(s)$ describing the late time relaxation of KPZ fluctuations with infinite boundary slopes. Additional checks of the main results against numerical simulations of TASEP are also provided.
@@ 2 Generator for the current
3 Perron-Frobenius stationary eigenstate
@@ II Iterative construction of the stationary eigenstate
1 Matrix product representation
@@
2 Height representation
Following [20], we associate with the configurations $C$ and $C^{\prime}$ in (S32), (S33) the height functions $H_{i}$ and $H_{i}^{\prime}$, $i=0,\ldots,L$ as in (5) and consider the path $P_{i}=H_{i}^{\prime}-H_{i}$, such that $P_{0}=0$ and $P_{i+1}-P_{i}\in\{-1,0,1\}$ for all $i$. Then, we note that $P_{i}=P_{i-1}$ corresponds to $K_{n_{i}^{\prime},n_{i}}=1$ in (S32), (S33), $P_{i}=P_{i-1}+1$ to $K_{n_{i}^{\prime},n_{i}}=E$ and $P_{i}=P_{i-1}-1$ to $K_{n_{i}^{\prime},n_{i}}=D$. In order to compute $K_{n_{1}^{\prime},n_{1}}\ldots K_{n_{L}^{\prime},n_{L}}$ in (S32) and (S33), the constant portions of the path $P_{i}$ can then be erased, as well as the local minima using $DE=1$, see the graphical representation below for an example.
@@ 2 Height representation
with $N=\sum_{\ell=1}^{L}n_{\ell}$, $N^{\prime}=\sum_{\ell=1}^{L}n_{\ell}^{\prime}$ the total number of particles in the configurations $C$, $C^{\prime}$, and the maxima computed over all the sites $i=0,\ldots,L$. The exponent of $E$ (respectively $D$) is non-negative since $H_{0}=H_{0}^{\prime}=0$ (resp. $H_{L}=\frac{L}{2}-N$ and $H_{L}^{\prime}=\frac{L}{2}-N^{\prime}$). The matrix product representations (S32) and (S33) can finally be computed using the rest of the algebra (S34), leading for arbitrary $\alpha$, $\beta$ to (13) for $T(\gamma)$ and
@@
with $N^{1}$ and $N^{2}$ the total number of particles in the system for the configurations $C^{1}$ and $C^{2}$. We specialize from now on to the special case $\alpha=\beta=1$. Taking the limit $\alpha,\beta\to 1$ in (S36), $U(\gamma)$ reduces in particular to the form given below (13) in the letter, which is independent of $\gamma$ since all the non-zero matrix elements then have $\max(H^{2}-H^{1})=0$.
@@ III KPZ fluctuations
1 Stationary cumulants
The first stationary cumulants $c_{k}$, defined in (4), can be extracted from $\mu(s)=s+\chi(v)/2$, $s=\chi^{\prime}(v)$ by expanding $\chi$ given in (16) in powers of $\mathrm{e}^{v}$ [24]. One has
@@ 1 Stationary cumulants
2 Correction $\nu_{x}(s)$ to stationary large deviations
From (S30) and (1),
@@ 2 Correction $\nu_{x}(s)$ to stationary large deviations
[…]
@@
see section 2. Then, for deterministic initial condition associated with the discrete height function $H^{0}$, (6) leads to
@@
Additionally, from the conjectured exact expression (23) for $\omega_{\rm stat}$ obtained from Bethe ansatz numerics, the function $\nu_{0}(s)$ for stationary initial condition has the alternative explicit expression (26), see section d for a detailed derivation.
3 Initial conditions corresponding to macroscopic density profiles
@@
Additionally, from the conjectured exact expression (24) and below for empty and full initial condition, the function $\nu_{0}(s)$ is found to have the alternative explicit expression (25) for both narrow wedge initial conditions above, see section d for a detailed derivation.
4 Late time correction to the stationary cumulants of $h(x,\tau)$
The coefficients $d_{1}(x)$ and $d_{2}(x)$ in (4) can be extracted from the expansion in powers of $s$ of the expression (18) for $\mathrm{e}^{\nu_{0}(s)}$ with $n=2$. Several Brownian averages then need to be computed, see section 3. Using (S72), (S74) and the Brownian averages involving maxima in table 1 (and in some places the fact that the opposite of a Wiener process is also a Wiener process), one finds (19) for $d_{1}(x)$ and
@@ 4 Late time correction to the stationary cumulants of $h(x,\tau)$
Figure S4: Third and fourth cumulants of $h(0,\tau)$ plotted as a function of time for infinite boundary slopes and stationary initial condition. The data were obtained from simulations of open TASEP with $L=1023$ sites, boundary rates $\alpha=\beta=1$. The red, horizontal lines are the exact expression in (S51) and (S52) for $d_{3}(0)$ and $d_{4}(0)$. As seen in the inset, which represents simulations with $L=127,255,511$ (from top to bottom on the left, from bottom to top on the right), finite size corrections $\sim 1/L$ are present. The simulations were averaged over a number of independent realizations detailed in section 1.
Explicit expressions can be obtained for specific initial conditions. For stationary initial condition $h_{0}=h_{\rm st}$, using again the results in table 1, (19) and (4) give
@@
These results are in good agreement with simulations of TASEP, see figures 3 and S4.
@@
Figure S5: Two first cumulants of $h(0,\tau)$ plotted as a function of time for infinite boundary slopes and flat initial condition $h_{0}(x)=0$. The average (respectively variance) of $h(0,\tau)$ is displayed on the left (resp. right). The data were obtained from simulations of open TASEP with $L=4095$ (resp. $L=1023$) sites, prepared initially with $(L-1)/2$ particles located at the even sites, with boundary rates $\alpha=\beta=1$. The red, horizontal line is the exact expression (S53) (resp. the numerical value (S54)). As seen in the inset, which represents from bottom to top simulations with $L=255,1023,4095$ (resp. from top to bottom with $L=127,255,511,1023$), the small discrepancy observed at late times is consistent with a finite size correction $\sim 1/\sqrt{L}$ (resp. $\sim 1/L$). The simulations were averaged over a number of independent realizations detailed in section 1.
For flat initial condition $h_{0}=0$, one finds from (19) and (4)
@@
Figure S6: Average and third cumulant of the KPZ height $h(0,\tau)$, defined from TASEP as in (S46), with boundary rates $\alpha=\beta=1$, for empty (grey dots) and full (black dots) initial condition, plotted as a function of time $\tau$ as in (8). The red, horizontal line is the exact expression in (S55), (S57). Left: $\langle h(0,\tau)\rangle-c_{1}\,\tau$ with $L=4095$ (inset: $L=255,1023,4095$ from bottom to top, indicating a finite size correction $\sim L^{-1/2}$). Right: $\langle h(0,\tau)^{3}\rangle_{\rm c}-c_{3}\,\tau$ with $L=1023$. The simulations were averaged over a number of independent realizations detailed in section 1.
@@
Figure S7: Average of the KPZ height $\langle h(x,\tau)\rangle-c_{1}\tau$ at time $\tau=1\gg\tau_{*}$ plotted as a function of the position $x$, defined from TASEP with boundary rates $\alpha=\beta=1$ for stationary (left), empty (right, grey dots) and full (right, black dots) initial conditions. The dots correspond to simulations with $L=255,1023,4095$ (left: from top to bottom; right: from bottom to top), and indicate a finite size correction $\sim L^{-1/2}$. In each case, the red curve is the exact expression from (S59). The simulations were averaged over a number of independent realizations detailed in section 1.
@@
Figure S8: Shifted variance of the KPZ height $\mathrm{Var}(h(x,1))-\frac{\mathrm{Var}(h(0,1))+\mathrm{Var}(h(1,1))}{2}$ at time $\tau=1\gg\tau_{*}$ plotted as a function of the position $x$, defined from TASEP with $L=1023$ and boundary rates $\alpha=\beta=1$ for stationary (blue dots), flat (green; even sites occupied), empty (grey dots) and full (black dots) initial conditions. The simulations were averaged over a number of independent realizations detailed in section 1.

### v2

@@ Abstract
$\langle\mathrm{e}^{sh(x,\tau)}\rangle\simeq\mathrm{e}^{\tau\mu(s)+\nu_{x}(s)}\quad\text{when $\tau\to\infty$}\;.$ (3)
@@
A classical mapping of TASEP to an interface growth model, see figure 1, interprets empty (resp. occupied) sites as portions of interface going up (resp. down). Modulo a global shift of the interface, this defines a bijection between configurations of the particles on the lattice and discrete heights functions $H_{i}$, normalized with $H_{0}=0$, written from occupation numbers $n_{i}$ (with $n_{i}=1$ if site $i$ is occupied and $n_{i}=0$ if site $i$ is empty) as
@@
$i=0,\ldots,L$. Movements of particles on the lattice then induce deposition of square blocks on the local minima of the interface, with $H_{i}\to H_{i}+1$ when a particle moves from site $i$ to site $i+1$ (with $i=0$ for particles entering at site $1$ and $i=L$ for particles leaving from site $L$). The resulting time dependent height function $H_{i}(t)$ is initialized at time $t=0$ as (5) with the occupation numbers of the initial configuration (which may be random). Additionally, the time-integrated current $Q_{i}(t)$, defined as the total number of particles that have moved from site $i$ to $i+1$ up to time $t$, is equal to
@@
Incidentally, the identity $H_{i}(t)=Q_{0}(t)+\sum_{\ell=1}^{i}\big{(}\frac{1}{2}-n_{\ell}(t)\big{)}$ holds, and there is thus a proper bijection at any time $t$ between height functions $H_{i}(t)$ and the joint data of the current at a given site and the configuration of particles.
@@
In the whole maximal current phase $\alpha>1/2$, $\beta>1/2$ of TASEP [27, 28, 29], the large scale limit $L\to\infty$ of the discrete height on the time scale $t\sim L^{3/2}$ leads to the KPZ fixed point with infinite boundary slopes, and we set already $\alpha=\beta=1$ for simplicity in the following. Considering a smooth enough continuous function $h_{0}(x)$, $0\leq x\leq 1$, with $h_{0}(0)=0$, we choose an initial condition of the form
@@
The matrix elements of the generator $M_{i}(\gamma)$ are equal to those of the Markov operator of TASEP, multiplied by $\mathrm{e}^{\gamma}$ for all transitions where a particle moves from site $i$ to site $i+1$. One has in particular $M_{i}(\gamma)=V_{i}^{-1}(\gamma)M_{0}(\gamma)V_{i}(\gamma)$ with $V_{i}$ diagonal such that $\langle C|V_{i}|C\rangle=\mathrm{e}^{-\gamma H_{i}}$, and $H_{i}$ related to the configuration $C$ of particles on the lattice through (5). The row vector $\langle\Sigma|=(1,1,\ldots,1)$ in (10) leads to a summation over all possible final states, and $|\mathbb{P}_{0}\rangle$ is the column vector of initial probabilities.
@@
where $w_{1}(u)$ and $w_{2}(u)$ are associated with $C^{1}$ and $C^{2}$, and the minimum is taken over all $u\in[0,1]$. Similarly, $\langle C^{2}|U|C^{1}\rangle$ leads to conditioning the corresponding $w_{1}$ and $w_{2}$ on $w_{1}(u)\geq w_{2}(u)$ for $u\in[0,1]$ and $w_{1}(1)=w_{2}(1)$. Equivalently, $w(x)=(w_{1}(x)+w_{2}(x))/\sqrt{2}$ and $e(x)=(w_{1}(x)-w_{2}(x))/\sqrt{2}$ are respectively a Wiener process and a Brownian excursion, independent from each other, see the Supplementary Material [34]. Then, (11) combined with (6), (Approach to stationarity for the KPZ fixed point with boundaries), (13) and (14) finally leads to our first main result, valid for any $n\in\mathbb{N}$ and arbitrary initial condition $h_{0}$
@@
The average height $\langle h(x,\tau)\rangle$ has the form (4) with the known value $c_{1}=3/2$. From (18), the new correction to stationarity (see the Supplementary Material [34] for the computation of the required Brownian averages) is
@@
$y_{0}^{+}$$y_{0}^{-}$$y_{1}^{+}$$y_{1}^{-}$$y_{2}^{+}$$y_{2}^{-}$$y_{3}^{+}$$y_{3}^{-}$$y_{4}^{+}$$y_{4}^{-}$
Figure 4: Image of the slit plane $\mathbb{C}\setminus[0,\infty)$ by the Bethe root functions $y_{j}^{\pm}(B)$ for open TASEP with $L=3$ and $\alpha=\beta=1$.
The $L+2$ solutions $w_{j}$, $j=0,\ldots,L+1$ of (22) lead to $y_{j}=y_{j}^{\pm}=w_{j}\pm w_{j}\sqrt{1-w_{j}^{-2}}$. Taking the principal branch for the square root, we note that the image of the branch cut $w_{j}\in(-\infty,-1]\cup[1,\infty)$ is the unit circle $|y_{j}^{\pm}|=1$, and one has $|y_{j}^{+}|>1$, $|y_{j}^{-}|<1$ outside the cut, see figure 4. Our numerics indicate that any eigenstate of $M_{i}(\gamma)$ can be generated by choosing Bethe roots $y_{j}=y_{j}^{\theta_{j}}$ with the signs $\theta_{j}=\pm$ such that
$\prod_{j=0}^{L+1}\theta_{j}=(-1)^{L}\;.$ (23)
The stationary eigenstate corresponds in particular to selecting $y_{j}=y_{j}^{-}$ for all $j$.
As a function of the parameter $B$, the Bethe roots $y_{j}^{\pm}$ are analytic outside the interval $[0,\infty)$, with the three branch points 00, $B_{*}>0$, $\infty$, where
$B_{*}=\frac{(L+2)^{L+2}}{2^{2L+4}(L+1)^{L+1}}\;.$ (24)
Analytic continuation across the two branch cuts $(0,B_{*})$ and $(B_{*},\infty)$ connects the various eigenstates of $M_{i}(\gamma)$ on the spectral curve $\det(\lambda\,\mathrm{Id}-M_{i}(\gamma))=0$. Labelling the $y_{j}^{\pm}$ as in figure 4, we observe that analytic continuation across the cuts $(B_{*},\infty)$ and $(0,B_{*})$ is realized by the action on the set $\{j_{\theta_{j}},j=0,\ldots,L+1\}$ of the respective permutations (written as a product of two cycles) $\mathfrak{a}=(0_{+},0_{-})\,(1_{+},2_{+},\ldots,(L+1)_{+},1_{-},2_{-},\ldots,(L+1)_{-})$ and $\mathfrak{b}=(0_{+},1_{+},\ldots,(L+1)_{+})\,(0_{-},1_{-},\ldots,(L+1)_{-})$. The condition (23) for acceptable choices of the Bethe roots $y_{j}=y_{j}^{\theta_{j}}$ is in particular preserved by $\mathfrak{a}$ and $\mathfrak{b}$, which gives an independent check of its validity.
@@
and $\omega^{\rm full}=\omega^{\rm empty}-\frac{L}{2}\,\kappa\,\frac{\mathrm{d}B}{B}$. These conjectures were checked against numerical diagonalization of $M_{0}(\gamma)$ up to system size $L=7$, with $100$ digits of precision, for a few generic values of $B$, and for all acceptable choices of signs $\theta_{j}$ satisfying (23).
When $L\to\infty$ under KPZ scaling (15) for $\gamma$, setting $B/B_{*}=-\mathrm{e}^{v-1}$, with $B_{*}$ in (24), a residue computation leveraging the fact that only the $y_{j}^{-}$ vanish when $B\to 0$ leads for the stationary eigenstate to $\kappa_{\rm st}\simeq 2\chi^{\prime\prime}(v)/\sqrt{L}$ with $\chi$ as in (16) and $v<1$ the unique solution of $\chi^{\prime}(v)=s$ for $s\in(0,s_{*}]$ with $s_{*}=\chi^{\prime}(1)\approx 0.145931$. Noting that $\mathcal{N}(B)$ evaluated at the stationary eigenstate appears in (11), we finally obtain our second main result for infinite boundary slopes,
@@
for narrow wedge initial condition at the left boundary, using $\omega^{\rm empty}$ in (26). The same result is obtained for a narrow wedge at the right boundary, using $\omega^{\rm full}$, with an additional deterministic subtraction in (9), see the Supplementary Material [34]. For stationary initial condition $h_{0}=h_{\rm st}$ defined in (2), one finds instead from (25)
@@
Supplementary Material
We give in this supplementary material a more detailed derivation of the main results presented in the letter for the function $\nu_{x}(s)$ describing the late time relaxation of KPZ fluctuations with infinite boundary slopes. Additional checks of the main results against numerical simulations of TASEP are also provided.
@@ I.2 Generator for the current
I.3 Perron-Frobenius stationary eigenstate
@@ II Iterative construction of the stationary eigenstate
II.1 Matrix product representation
@@
II.2 Height representation
Following [20], we associate with the configurations $C$ and $C^{\prime}$ in (S34), (S35) the height functions $H_{i}$ and $H_{i}^{\prime}$, $i=0,\ldots,L$ as
$H_{i}=\sum_{\ell=1}^{i}\Big{(}\frac{1}{2}-n_{\ell}\Big{)}\qquad\text{and}\qquad H_{i}^{\prime}=\sum_{\ell=1}^{i}\Big{(}\frac{1}{2}-n_{\ell}^{\prime}\Big{)}$ (S37)
with $n_{i}$, $n_{i}^{\prime}$ the occupation numbers of the configurations $C$ and $C^{\prime}$, and consider the path $P_{i}=H_{i}^{\prime}-H_{i}$, such that $P_{0}=0$ and $P_{i+1}-P_{i}\in\{-1,0,1\}$ for all $i$. Then, we note that $P_{i}=P_{i-1}$ corresponds to $K_{n_{i}^{\prime},n_{i}}=1$ in (S34), (S35), $P_{i}=P_{i-1}+1$ to $K_{n_{i}^{\prime},n_{i}}=E$ and $P_{i}=P_{i-1}-1$ to $K_{n_{i}^{\prime},n_{i}}=D$. In order to compute $K_{n_{1}^{\prime},n_{1}}\ldots K_{n_{L}^{\prime},n_{L}}$ in (S34) and (S35), the constant portions of the path $P_{i}$ can then be erased, as well as the local minima using $DE=1$, see the graphical representation below for an example.
@@ II.2 Height representation
with $N=\sum_{\ell=1}^{L}n_{\ell}$, $N^{\prime}=\sum_{\ell=1}^{L}n_{\ell}^{\prime}$ the total number of particles in the configurations $C$, $C^{\prime}$, and the maxima computed over all the sites $i=0,\ldots,L$. The exponent of $E$ (respectively $D$) is non-negative since $H_{0}=H_{0}^{\prime}=0$ (resp. $H_{L}=\frac{L}{2}-N$ and $H_{L}^{\prime}=\frac{L}{2}-N^{\prime}$). The matrix product representations (S34) and (S35) can finally be computed using the rest of the algebra (S36), leading for arbitrary $\alpha$, $\beta$ to
$\langle C^{2}|T(\gamma)|C^{1}\rangle=\mathrm{e}^{-\gamma\,\max(H^{2}-H^{1})}\;,$ (S39)
and
@@
with $N^{1}$ and $N^{2}$ the total number of particles in the system for the configurations $C^{1}$ and $C^{2}$. We specialize from now on to the special case $\alpha=\beta=1$. Taking the limit $\alpha,\beta\to 1$ in (S40), $U(\gamma)$ reduces in particular to
[…]
All the non-zero matrix elements $\langle C^{2}|U(\gamma)|C^{1}\rangle$ have $\max(H^{2}-H^{1})=0$, which explains why $U(\gamma)$ is then independent of $\gamma$.
@@ III KPZ fluctuations
III.1 Stationary cumulants
The first stationary cumulants $c_{k}$, defined in (4), can be extracted from $\mu(s)=s+\chi(v)/2$, $s=\chi^{\prime}(v)$ by expanding
$\chi(v)=-\frac{1}{4\sqrt{\pi}}\sum_{k=1}^{\infty}\frac{(2k)!}{k^{k+5/2}\,k!}\,\Big{(}-\frac{\mathrm{e}^{v}}{4}\Big{)}^{k}\;.$ (S42)
in powers of $\mathrm{e}^{v}$ [24]. One has
@@ III.1 Stationary cumulants
III.2 Correction $\nu_{x}(s)$ to stationary large deviations
From (S32) and (II.1),
@@ III.2 Correction $\nu_{x}(s)$ to stationary large deviations
The matrix elements of $T(\gamma)$ are given in (S39) in terms of the corresponding discrete height functions from (S37). At large $L$, the KPZ fixed point with infinite boundary slopes is expected in the whole maximal current phase $\alpha,\beta>1/2$ (fine-tuning $\alpha-1/2\sim\beta-1/2\sim L^{-1/2}$ would be required for finite boundary slopes). For simplicity, we set $\alpha=\beta=1$, so that the matrix elements of $U(\gamma)$ are given by (S41). Then, Donsker’s theorem (14) leads at large $L$ with $i=xL$, $\gamma$ scaling as in (15) and an initial condition of the form (7) to the following Brownian average
@@
see section IV.2. Then, for deterministic initial condition associated with the discrete height function $H^{0}$, (6) leads to
@@
Additionally, from the conjectured exact expression (25) for $\omega_{\rm stat}$ obtained from Bethe ansatz numerics, the function $\nu_{0}(s)$ for stationary initial condition has the alternative explicit expression (28), see section V.2.d for a detailed derivation.
III.3 Initial conditions corresponding to macroscopic density profiles
@@
Additionally, from the conjectured exact expression (26) and below for empty and full initial condition, the function $\nu_{0}(s)$ is found to have the alternative explicit expression (27) for both narrow wedge initial conditions above, see section V.2.d for a detailed derivation.
III.4 Late time correction to the stationary cumulants of $h(x,\tau)$
The coefficients $d_{1}(x)$ and $d_{2}(x)$ in (4) can be extracted from the expansion in powers of $s$ of the expression (18) for $\mathrm{e}^{\nu_{0}(s)}$ with $n=2$. Several Brownian averages then need to be computed, see section IV.3. Using (S78), (S80) and the Brownian averages involving maxima in table 1 (and in some places the fact that the opposite of a Wiener process is also a Wiener process), one finds (19) for $d_{1}(x)$ and
@@ III.4 Late time correction to the stationary cumulants of $h(x,\tau)$
Figure S5: Third and fourth cumulants of $h(0,\tau)$ plotted as a function of time for infinite boundary slopes and stationary initial condition. The data were obtained from simulations of open TASEP with $L=1023$ sites, boundary rates $\alpha=\beta=1$. The red, horizontal lines are the exact expression in (S57) and (S58) for $d_{3}(0)$ and $d_{4}(0)$. As seen in the inset, which represents simulations with $L=127,255,511$ (from top to bottom on the left, from bottom to top on the right), finite size corrections $\sim 1/L$ are present. The simulations were averaged over a number of independent realizations detailed in section I.1.
Explicit expressions can be obtained for specific initial conditions. For stationary initial condition $h_{0}=h_{\rm st}$, using again the results in table 1, (19) and (III.4) give
@@
These results are in good agreement with simulations of TASEP, see figures 3 and S5.
@@
Figure S6: Two first cumulants of $h(0,\tau)$ plotted as a function of time for infinite boundary slopes and flat initial condition $h_{0}(x)=0$. The average (respectively variance) of $h(0,\tau)$ is displayed on the left (resp. right). The data were obtained from simulations of open TASEP with $L=4095$ (resp. $L=1023$) sites, prepared initially with $(L-1)/2$ particles located at the even sites, with boundary rates $\alpha=\beta=1$. The red, horizontal line is the exact expression (S59) (resp. the numerical value (S60)). As seen in the inset, which represents from bottom to top simulations with $L=255,1023,4095$ (resp. from top to bottom with $L=127,255,511,1023$), the small discrepancy observed at late times is consistent with a finite size correction $\sim 1/\sqrt{L}$ (resp. $\sim 1/L$). The simulations were averaged over a number of independent realizations detailed in section I.1.
For flat initial condition $h_{0}=0$, one finds from (19) and (III.4)
@@
Figure S7: Average and third cumulant of the KPZ height $h(0,\tau)$, defined from TASEP as in (S52), with boundary rates $\alpha=\beta=1$, for empty (grey dots) and full (black dots) initial condition, plotted as a function of time $\tau$ as in (8). The red, horizontal line is the exact expression in (S61), (S63). Left: $\langle h(0,\tau)\rangle-c_{1}\,\tau$ with $L=4095$ (inset: $L=255,1023,4095$ from bottom to top, indicating a finite size correction $\sim L^{-1/2}$). Right: $\langle h(0,\tau)^{3}\rangle_{\rm c}-c_{3}\,\tau$ with $L=1023$. The simulations were averaged over a number of independent realizations detailed in section I.1.
@@
Figure S8: Average of the KPZ height $\langle h(x,\tau)\rangle-c_{1}\tau$ at time $\tau=1\gg\tau_{*}$ plotted as a function of the position $x$, defined from TASEP with boundary rates $\alpha=\beta=1$ for stationary (left), empty (right, grey dots) and full (right, black dots) initial conditions. The dots correspond to simulations with $L=255,1023,4095$ (left: from top to bottom; right: from bottom to top), and indicate a finite size correction $\sim L^{-1/2}$. In each case, the red curve is the exact expression from (S65). The simulations were averaged over a number of independent realizations detailed in section I.1.
@@
Figure S9: Shifted variance of the KPZ height $\mathrm{Var}(h(x,1))-\frac{\mathrm{Var}(h(0,1))+\mathrm{Var}(h(1,1))}{2}$ at time $\tau=1\gg\tau_{*}$ plotted as a function of the position $x$, defined from TASEP with $L=1023$ and boundary rates $\alpha=\beta=1$ for stationary (blue dots), flat (green; even sites occupied), empty (grey dots) and full (black dots) initial conditions. The simulations were averaged over a number of independent realizations detailed in section I.1.